OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y = 4x - 11 y = 4x + 11 y = 4x + 7 y = 4x - 7

OpenStudy (anonymous):

WILL FAN AND MEDAL

OpenStudy (anonymous):

Plug in the -3 for x and -1 for y since the slope is the same in all equations - can not eliminate up front

OpenStudy (anonymous):

-1 = 4(-3) + m -1 = -12 + m -1 + 12 = m 11 = m which tells me the y - intercept is 11 - so therefore B is the correct answer

OpenStudy (anonymous):

thanks but how did you get the y-intercept (my first year of online homeschooling and it's not very easy)

OpenStudy (anonymous):

I just solved for B (I used M but should of used B) note Y = mx + b M = slope B = y-intercept so if you are give Y, M, and X solve for B